CAN SOMEONE HELP ME PLEASE this problem is confusing me. i don't get how to do it.
triangle RST has vertices R(3,2),S(-1,6), and T(-5,-6).
Use the Pythagorean Theorem to show that triangle RST is a right triangle

Answer by praseenakos@yahoo.com(507)



R(3,2)=(X1,Y1);S(-1,6)=(X2,Y2); T(-5,-6)=(X3,Y3)
use distance formula to find RS,ST,&TR
RS squared = (X2-x1)squared+(Y2-Y1)squared= (-4)squared+(4)squared=16+16=32
ST squared=(X3-X2)squared+(Y3-Y2)squared= (-4)squared+(-12)squared=16+144=160
TR squared=(X3-X1)squared+(Y3-Y1)squared= (-8)squared+(-8)squared=64+64=128
from the above we can conclude that RS squared+TR squared = ST squared
therefore RST is a right triangle.